Divers in Acapulco fall 36 m from a cliff into the water. Estimate the time interval needed to reach the water Express the answer to three significant figures and include the appropriate units. v = 26.6 m/s.
Answer #1

From kinematic equation

s=u*t+1/2*a*t^2

at rest, u=0

h=1/2*g*t^2

36=1/2*9.8*t^2

t=sqrt((36*2/9.8))

t=2.71 sec

the time interval needed to reach the water is 2.71 sec

---------------------------------------------------------------------

From kinematic equation

v=u+a*t

v=0+g*t

v=9.8*2.71

v =26.6 m/sec

speed when they enter the water is 26.6 m/sec